a.) Find Laplace transform F(s) of (3-e2t + 5 e 6t) sin(7+)
Answer #1

Ques Soln to. Given Flev = (3-c** *524) sin(78) 08 Elt= 35.n(76) e2+ sin(at) tse 6+ Sin (74) we will find Laplace-transform o= ) Sin (7+1 = it 2 -fit zi Soi F(S) = I zi ** (*** ***) (3-2-se). = F(S) = -(S-7i)+ -(S+70+ - e 2i slere 2013 (3-2*+s=/ F7 F(S)= - 3 -CS-1 ) 2 3. (s-) + exstaigt (s :) + -(s--t S+71-2)t (s-2-1) (s-24) S (Sc) 4 ਏ S (3-4) -(S+7i+6) + ( 3 44 4k | =)=) FIS) = 20 + (stril S- 2470 S + s (S+6-72) (5+6+71) = F (s) = 3 20 [ (s+: -5+76) (52+72) + (5-2-77-542tti) (5-2)²+72 + S S+